# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v1

@@ Abstract
The Chiral Magnetic Effect (CME) is a phenomenon in which electric charge is separated by a strong magnetic field from local domains of chirality imbalance and parity violation in quantum chromodynamics (QCD). The CME-sensitive observable, charge-dependent three-point azimuthal correlator $\Delta\gamma$, is contaminated by a major physics background proportional to the particle’s elliptic anisotropy $v_{2}$. Event-shape engineering (ESE) binning events in dynamical fluctuations of $v_{2}$ and event-shape selection (ESS) binning events in statistical fluctuations of $v_{2}$ are two methods to search for the CME by projecting $\Delta\gamma$ to the $v_{2}=0$ intercept. We conduct a systematic study of these two methods using physics models as well as toy model simulations. It is observed that the ESE method requires significantly more statistics than the ESS method to achieve the same statistical precision of the intercept. It is found that the intercept from the ESS method depends on the details of the event content, such as the mixtures of background-contributing sources, and thus is not a clean measure of the CME.
@@ I Introduction
It has been predicted by quantum chromodynamics (QCD) that vacuum fluctuations can result in gluon fields of non-zero topological charges in local metastable domains. Interactions of quarks with such gluon fields can cause chirality imbalance, breaking the parity and charge-parity symmetries in those domains. Such chirality imbalance, under a strong magnetic field, would result in charge separation, a phenomenon called the chrial magnetic effect (CME) [1, 2, 3, 4].
Ultra-strong magnetic fields are presumably produced in non-head-on relativistic heavy ion collisions [5, 6]. The magnetic field is on average perpendicular to the reaction plane (RP, the plane span by the beam and the impact parameter direction of the collision). The CME-induced charge separation is along the direction of the magnetic field, and may be conveniently quantified by the $a_{1}$ variable in Fourier series of particle azimuthal distributions [7],
$dN_{\pm}/d\phi^{*}\propto 1\pm 2a_{1}\sin\phi^{*}+2v_{2}\cos 2\phi^{*}+\cdots\,,$ (1)
where $\phi^{*}$ is the azimuthal angle of the particle momentum vector with respect to the RP and the subscript ‘$\pm$’ indicates particle charge sign. The elliptic anisotropy $v_{2}$ harmonic is the leading modulation in particle distributions produced in relativistic heavy ion collisions [8]. Because of a vanishing mean $a_{1}$ due to random fluctuations, a commonly used observable is the three-point correlator [7],
@@
where $\phi_{\alpha}$ and $\phi_{\beta}$ are the azimuthal angles of two particles of interest (POI). To cancel charge-independent backgrounds, such as effects from global momentum conservation, the difference between opposite-sign (OS) and same-sign (SS) correlators is used [9],
@@
Definite signals of the CME have not yet been observed [10, 11, 12, 9, 13]. The major difficulty is the large contamination in $\Delta\gamma$ from mandune QCD backgrounds. Those backgrounds arise from genuine two-particle correlations, such as correlations between daughter particles from a resonance decay, or among particles from the same jet or back-to-back dijet, coupled with elliptic anisotropies of those background sources [7, 14, 15, 16]. This background correlation can be schematically expressed as [7, 17]
@@
Large efforts have since been invested to eliminate or mitigate those backgrounds [18, 19, 20, 21, 22], including innovative observables [23, 24, 25, 26, 27, 28]. One of the techniques is the event-shape engineering (ESE) [29, 23, 24] method, grouping events into classes of different $v_{2}$ values relying on the dynamical fluctuations of $v_{2}$. A variation of this method is to select events according to the particle emission pattern, called the event-shape selection (ESS) method [30], relying on the statistical fluctuations of $v_{2}$. The main difference between the two methods is in the ways of quantifying the event shape. The ESE method [29] uses a variable calculated not from the particles of interest (POI) that are used for the physics measurement of $\Delta\gamma$, while the ESS method uses a variable calculated from the POIs [30]. Each method analyzes $\Delta\gamma$ as a function of $v_{2}$ in events binned according to the corresponding event-shape variable and projects $\Delta\gamma$ to $v_{2}=0$ to obtain the intercept, presumably insensitive to backgrounds. While the ESE method is relatively straightforward to comprehend, the ESS method is complex because the POIs are used in both event selection and physics measurement. In this paper, we conduct a systematic study of the ESS and ESE methods for the CME search using several physics models as well as toy models of varying ingredients, with the goal to elucidate what the projected intercept entails from each method.
The rest of the paper is organized as follows. Section II describes the details of the ESE and ESS methods (and a related early method [18]) used for the CME search. Section III gives brief descriptions of the models used in this study. Section IV presents our model simulation results using ESS and ESE and discusses the findings. Finally, a summary is given in Section V. The appendix compiles all the $\Delta\gamma$ vs. $v_{2}$ plots from the simulations by the ESS and ESE methods.
@@ II The Event-Shape Techniques
where $N$ is the number of particles in the event within a given phase space (we use “particles” in our description for convenience without loss of generality). One may use the normalized $\hat{q}_{2}$ quantity [30] for ESE,
@@
In ESE, the $q_{2}^{2}$ is typically measured at forward and backward rapidities [23, 24]. The event classes selected by $q_{2}$ (or equivalently $q_{2}^{2}$ or $\hat{q}_{2}^{2}$) within a given narrow centrality range will have different average $\left\langle v_{2}\right\rangle$ in the midrapidity region, or any other phase space regions displaced from that for $q_{2}^{2}$ computation. In other words, because $q_{2}$ is correlated with $\left\langle v_{2}\right\rangle$, ESE is dividing events in groups of varying $\left\langle v_{2}\right\rangle$, which dynamically fluctuates from event to event. These dynamical fluctuations are mainly caused by initial geometry fluctuations within the given narrow centrality bin [31, 32].
@@
Figure 1: Cartoons illustrating the various event-shape methods in terms of the event-by-event resonance $v_{\rm 2,{\rm res}}$ vs. final-state single particle $v_{2}$ in events within a narrow centrality bin. Here $v_{\rm 2,{\rm res}}$ and $v_{2}$ are taken to be the event-by-event elliptic shape variables, $\left\langle\cos 2(\phi-\psi)\right\rangle$, with respect to the event plane $\psi$ from another phase space different from the particles of interest (POI). (Left panel) Event-Shape Engineering (ESE) where events are selected according to $q_{2}$ in phase space region displaced from that of POIs’, in which the average $\left\langle v_{\rm 2,\rho}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$, unbiased, and the spread in $\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ are due to dynamical fluctuations. (Center penal) Event-Shape (ES) method where events are binned in the observed elliptic shape variable $v_{2}^{\rm obs}$ of POIs, and the average $v_{\rm 2,{\rm res}}$ in those events are strongly biased, not strictly proportional to $v_{2}^{\rm obs}$ and with a positive $v_{\rm 2,{\rm res}}$ value at $v_{2}^{\rm obs}=0$. The wide range in $v_{2}^{\rm obs}$ (including negative values) is mainly due to statistical fluctuations. (Right panel) Event-Shape Selection (ESS) where events are selected according to $\hat{q}_{2,{\rm pair}}^{2}$ of pairs of POIs in which the $v_{\rm 2,{\rm res}}$ is unnecessarily proportional to $v_{2}$ due to complicated biases from the $\hat{q}_{2,{\rm pair}}^{2}$ selection using the same POIs as for $v_{2}$. The wide spreads in $v_{2}$ and $v_{\rm 2,{\rm res}}$ are mainly due to statistical fluctuations inherited in the ESS method involving the same POIs.
@@
Since the CME background comes from the elliptic anisotropies imprinted in the particles of interests (POI), it is interesting to select events with an event-shape ellipticity $v_{2}^{\rm obs}\equiv\left\langle\cos 2(\phi-\psi)\right\rangle$ calculated by POIs themselves (where the EP $\psi$ is determined in another phase space away from the POIs) and examine the $\Delta\gamma$ observable as a function of $v_{2}^{\rm obs}$. This method selects on the event-by-event elliptic-shape quantity (particle emissin pattern) $v_{2}^{\rm obs}$ which can fluctuate even to negative values because of large statistical fluctuations, the intercept at $v_{2}^{\rm obs}=0$, more sensitive to the CME, can be well determined from data. This event-shape (ES) analysis was carried out by STAR [18] and the intercept of the linear fit is consistent with zero with the then-available statistics. However, the backgrounds are not determined by $v_{2}^{\rm obs}$ but by $\left\langle v_{\rm 2,{\rm res}}\right\rangle$. Statistically fluctuated $v_{2}^{\rm obs}=0$ does not necessarily guarantee $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$. This is illustrated by the cartoon in Fig. 1 (center). Although $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$ with an unbiased event selection, the strongly (maximally) biased selection of events by $v_{2}^{\rm obs}$ will obscure the relationship of the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and $v_{2}^{\rm obs}$ in each bin of $v_{2}^{\rm obs}$ of those events, as indicated by the hollow box in Fig. 1 (center). Because of the overall positive means $\left\langle v_{2}^{\rm obs}\right\rangle=\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ over the entire event sample, the linear projection will give a positive intercept of $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ at $v_{2}^{\rm obs}=0$. In other words, with finite positive average $\left\langle v_{2}\right\rangle$ in heavy ion collisions, $\left\langle v_{\rm 2,{\rm res}}\right\rangle>0$ at $v_{2}^{\rm obs}=0$ with events binned in $v_{2}^{\rm obs}$. This has been verified by model study [34]. Thus, positive residual background remains in the linear fit intercept in the analysis of $\Delta\gamma$ as a function of $v_{2}^{\rm obs}$ [18], although the background is significantly reduced compared to that in the inclusive $\Delta\gamma$ measurement. Ideally, in the ES analysis, one would want to select events with the $v_{2}^{\rm obs}$ values of all resonances (background sources) to be zero, however, such selection is experimentally insurmountable.
@@
It has been recently proposed by Xu et al. [30] to select events according to the $\hat{q}_{2,\textsc{poi}}^{2}$ of the POIs, and study the $\Delta\gamma$ as a function of $v_{2}$ of the POIs. This is similar in spirit to the ES method above where the events are selected by $v_{2}^{\rm obs}$ and the $\Delta\gamma$ is examined as a function of the same $v_{2}^{\rm obs}$ variable. Here, the selection variable $\hat{q}_{2,\textsc{poi}}^{2}$ and the $v_{2}$ variable are different, although related as they are computed by the same POIs. The $\hat{q}_{2}^{2}$ variable is an event-by-event quantity connected to two-particle cumumant anisotropy (cf Eq. 5), whereas $v_{2}$ can be measured with respect to $\psi$ from another phase space like in the ES method. (Of course $v_{2}$ can also be computed by the two-particle cumulant method, in which case it is more directly connected to $\hat{q}_{2}^{2}$.) The variations in the event-by-event $\hat{q}_{2,\textsc{poi}}^{2}$ quantity, within a given narrow centrality bin, are mainly from statistical fluctuations. Since the $v_{2}$ in those $\hat{q}_{2,\textsc{poi}}^{2}$-selected events are related to $\hat{q}_{2,\textsc{poi}}^{2}$ itself, the variations in $v_{2}$ is also primarily of statistical nature. (Note, we simply use $v_{2}$ here, instead of the $\left\langle v_{2}\right\rangle$ for ESE, to indicate that the $v_{2}$ in ESS is of statistical nature although it is an “average” $v_{2}$ in events selected by $\hat{q}_{2,\textsc{poi}}^{2}$.) Since this technique selects on the statistical fluctuations in event shape, not on the dynamical fluctuations in $\left\langle v_{2}\right\rangle$ like in the ESE method, it is called the Event-Shape-Selection (ESS) method [30]. The most important difference in the ESS method from the ESE method is that the $\hat{q}_{2}^{2}$ is computed from the POIs in the former but in the later the phase-space where to compute $\hat{q}_{2}^{2}$ is displaced from that of the POIs’. In both approaches, one groups events (within a given narrow centrality bin) according to $\hat{q}_{2}^{2}$ and studies the $\Delta\gamma$ variable as a function of the $v_{2}$ of the POIs.
@@
Several combinations of $\hat{q}_{2,\textsc{poi}}^{2}$ and $v_{2}$ variables have been investigated in Ref. [30]. The $q_{2,\textsc{poi}}^{2}$ can be computed using single particle azimuthal angles of the POIs by Eq. 5. It can also be defined by pairs of POIs similar to Eq. 5; in this case, the $q^{2}_{2,{\rm pair}}$ of pairs is calculated by substituting the $i^{\rm th}$ particle’s $\phi_{i}$ by the azimuthal angle of the $i^{\rm th}$ pair in Eq. 5, and the sum runs over all pairs of POIs in the event. The normalized $\hat{q}_{2,{\rm pair}}^{2}$ is then

### v3

@@ Abstract
The Chiral Magnetic Effect (CME) is a phenomenon in which electric charge is separated by a strong magnetic field from local domains of chirality imbalance and parity violation in quantum chromodynamics (QCD). The CME-sensitive observable, charge-dependent three-point azimuthal correlator $\Delta\gamma$, is contaminated by a major physics background proportional to the particle’s elliptic flow anisotropy $v_{2}$. Event-shape engineering (ESE) binning events in dynamical fluctuations of $v_{2}$ and event-shape selection (ESS) binning events in statistical fluctuations of $v_{2}$ are two methods to search for the CME by projecting $\Delta\gamma$ to the measured anisotropy $v_{2}=0$ intercept. We conduct a systematic study of these two methods using physics models as well as toy model simulations. It is observed that the ESE method fulfills the general premise of measuring the CME but is statistically hungry. It is found that the intercept from the ESS method depends on the details of the event content, such as the mixtures of background-contributing sources, because of statistical fluctuations of intertwining variables used in the method, and is thus not practically useful or clean to measure the CME.
@@ I Introduction
It has been predicted by quantum chromodynamics (QCD) that vacuum fluctuations can result in gluon fields of nonzero topological charges in local metastable domains. Interactions of quarks with such gluon fields can cause chirality imbalance, breaking the parity and charge-parity symmetries in those domains. Such chirality imbalance, under a strong magnetic field, would result in charge separation, a phenomenon called the chrial magnetic effect (CME) [1, 2, 3, 4].
Ultra-strong magnetic fields are presumably produced in non-head-on relativistic heavy ion collisions [5, 6]. The magnetic field is on average perpendicular to the reaction plane (RP, the plane span by the beam and the impact parameter direction of the collision). The CME-induced charge separation is along the direction of the magnetic field, and may be conveniently quantified by the $a_{1}$ variable in Fourier series of particle azimuthal distributions [7, 8],
$dN_{\pm}/d\tilde{\phi}\propto 1\pm 2a_{1}\sin\tilde{\phi}+2v_{2}\cos 2\tilde{\phi}+\cdots\,,$ (1)
where $\tilde{\phi}$ is the azimuthal angle of the particle momentum vector with respect to the RP and the subscript ‘$\pm$’ indicates particle charge sign. The elliptic anisotropy $v_{2}$ harmonic is the leading modulation in particle distributions produced in relativistic heavy ion collisions [9]. Because of a vanishing mean $a_{1}$ due to random fluctuations, a commonly used observable is the three-point correlator [8],
@@
where $\phi_{\alpha}$ and $\phi_{\beta}$ are the azimuthal angles of two particles of interest (POIs), and $\psi$ is that of the RP (or participant plane because of fluctuations in the collision overlap geometry [10]). To cancel charge-independent backgrounds, such as effects from global momentum conservation, the difference between opposite-sign (OS) and same-sign (SS) correlators is used [11, 12],
@@
Definite signals of the CME have not yet been observed [11, 12, 13, 14, 15, 16]. The major difficulty is the large contamination in $\Delta\gamma$ from mundane QCD backgrounds. Those backgrounds arise from genuine two-particle correlations, such as correlations between daughter particles from a resonance decay, or between particles from the same jet or back-to-back dijet, coupled with elliptic flow anisotropies of those background sources [8, 17, 18, 19, 20, 21]. This background correlation can be schematically expressed as [8, 16, 22, 23]
@@
Large efforts have since been invested to eliminate or mitigate these backgrounds [24, 25, 26, 27, 28], including innovative observables [29, 30, 31, 32, 33, 34]. One of the techniques is the event-shape engineering (ESE) [35, 29, 30] method, grouping events into classes of different $v_{2}$ values relying on the dynamical fluctuations of $v_{2}$. A variation of this method is to select events according to the particle emission pattern, called the event-shape selection (ESS) method [36], relying on the statistical fluctuations of $v_{2}$. The main difference between the two methods is in the ways of quantifying the event shape. The ESE method [35] uses a variable calculated not from the POIs that are used for the physics measurement of $\Delta\gamma$ and $v_{2}$, while the ESS method uses a variable calculated from the POIs [36]. Each method analyzes $\Delta\gamma$ as a function of the measured $v_{2}$ in events binned according to the corresponding event-shape variable and projects $\Delta\gamma$ to $v_{2}=0$ to obtain the intercept, presumably more sensitive to CME signals than the overall $\Delta\gamma$ measurement and less sensitive (or ideally insensitive) to backgrounds.
The ESE method is relatively straightforward to comprehend. The ESS method is, however, complex because the POIs are used in both event selection and physics measurements. In this paper, we conduct a systematic study of the ESS and ESE methods for the CME search using several physics models as well as toy models of varying ingredients, with the goal to elucidate what the projected intercept entails from each method.
The focus of our study is placed on the effects of the event classifying variables on the intercept under various circumstances of the simulated events. We note, however, that the $v_{2}$ measured in experiments contains not only the flow anisotropy but also the so-called nonflow correlations, dependent of analysis methods. Nonflow correlations refer to those unrelated to the collision geometry common to the entire event, such as resonance decays and (di-)jet correlations [37, 38, 39, 40]. While we touch upon the nonflow issue in this article, a full and thorough discussion of nonflow effects on the ESE and ESS methods is outside the scope of the present study.
The rest of the paper is organized as follows. Section II describes the details of the ESE and ESS methods (and a related early method [24]) used for the CME search. Section III gives brief descriptions of the models used in this study. Section IV presents our model simulation results using ESS and ESE and discusses the findings. Finally, a summary is given in Section V. The appendix compiles all the $\Delta\gamma$ vs. $v_{2}$ plots from the simulations using the ESS and ESE methods.
@@ II The Event-Shape Techniques
where $N$ is the number of particles in the event within a given kinematic region (we use “particles” in our description for convenience without loss of generality). One may use the normalized $\hat{q}_{2}$ quantity [36] for ESE,
@@
In ESE, the $q_{2}^{2}$ is typically measured at forward and backward rapidities [29, 30]. The event classes selected by $q_{2}$ (or equivalently $q_{2}^{2}$ or $\hat{q}_{2}^{2}$) within a given narrow centrality range will have a different average $\left\langle v_{2}\right\rangle$ in the midrapidity region, or any other kinematic regions displaced from that for $q_{2}^{2}$ computation. In other words, because $q_{2}$ is correlated with $\left\langle v_{2}\right\rangle$, ESE is dividing events in groups of varying $\left\langle v_{2}\right\rangle$, which dynamically fluctuates from event to event. These dynamical fluctuations are mainly caused by initial geometry fluctuations within the given narrow centrality bin [41, 42].
@@
Figure 1: Cartoons illustrating the various event-shape methods in terms of the event-by-event resonance $v_{\rm 2,{\rm res}}$ vs. final-state single particle $v_{2}$ in events within a narrow centrality bin. Here $v_{\rm 2,{\rm res}}$ and $v_{2}$ are taken to be the event-by-event elliptic shape variables, $\left\langle\cos 2(\phi-\psi)\right\rangle$, with respect to the event plane $\psi$ from another kinematic region different from the particles of interest (POIs). (Left panel) Event-Shape Engineering (ESE) where events are selected according to $q_{2}$ in kinematic region displaced from that of POIs’ (each $q_{2}$-class is indicated by the brown ellipse), in which the average $\left\langle v_{\rm 2,\rho}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$, unbiased, and the spread in $\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ (indicated by the red circles) are due to dynamical fluctuations. (Center penal) Event-Shape (ES) method where events are binned in the observed elliptic shape variable $v_{2}^{\rm obs}$ of POIs (one bin is indicated by the brown rectangle), and the average $v_{\rm 2,{\rm res}}$ in those events (indicated by the red circles) are strongly biased, not strictly proportional to $v_{2}^{\rm obs}$ and with a positive $v_{\rm 2,{\rm res}}$ value at $v_{2}^{\rm obs}=0$. The wide range in $v_{2}^{\rm obs}$ (including negative values) is mainly due to statistical fluctuations. (Right panel) Event-Shape Selection (ESS) where events are selected according to $\hat{q}_{2,{\rm pair}}^{2}$ of pairs of POIs (each $\hat{q}_{2,{\rm pair}}^{2}$-class is indicated by the brown ellipse) in which the $v_{\rm 2,{\rm res}}$ is unnecessarily proportional to $v_{2}$ (indicated by the red circles) due to complicated biases from the $\hat{q}_{2,{\rm pair}}^{2}$ selection using the same POIs as for $v_{2}$. The wide spreads in $v_{2}$ and $v_{\rm 2,{\rm res}}$ are mainly due to statistical fluctuations inherited in the ESS method involving the same POIs. The green cross in each cartoon indicates the overall averages $\left\langle\left\langle v_{2}\right\rangle\right\rangle$ and $\left\langle\left\langle v_{\rm 2,{\rm res}}\right\rangle\right\rangle$ over the entire event sample.
@@
Since the CME background comes from the elliptic anisotropies imprinted in the POIs, it is interesting to select events with an event-shape ellipticity $v_{2}^{\rm obs}\equiv\left\langle\cos 2(\phi-\psi)\right\rangle$ calculated by POIs themselves (where the EP $\psi$ is determined in another kinematic region away from the POIs) and examine the $\Delta\gamma$ observable as a function of $v_{2}^{\rm obs}$. This method selects on the event-by-event elliptic-shape quantity (particle emissin pattern) $v_{2}^{\rm obs}$ which can fluctuate even to negative values because of large statistical fluctuations, the intercept at $v_{2}^{\rm obs}=0$, more sensitive to the CME, can be well determined from data. This event-shape (ES) analysis was performed by STAR [24] and the intercept of the linear fit is consistent with zero with the then-available statistics. However, the backgrounds are not determined by $v_{2}^{\rm obs}$ but by $\left\langle v_{\rm 2,{\rm res}}\right\rangle$. Statistically fluctuated $v_{2}^{\rm obs}=0$ does not necessarily guarantee $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$. This is illustrated by the cartoon in Fig. 1 (center). Although $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$ with an unbiased event selection, the strongly (maximally) biased selection of events by $v_{2}^{\rm obs}$ will obscure the relationship of the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and $v_{2}^{\rm obs}$ in each bin of $v_{2}^{\rm obs}$ of those events, as indicated by the hollow box in Fig. 1 (center). Because of the overall positive means $\left\langle v_{2}^{\rm obs}\right\rangle=\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ over the entire event sample, the linear projection will give a positive intercept of $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ at $v_{2}^{\rm obs}=0$. In other words, with finite positive average $\left\langle v_{2}\right\rangle$ in heavy ion collisions, $\left\langle v_{\rm 2,{\rm res}}\right\rangle>0$ at $v_{2}^{\rm obs}=0$ with events binned in $v_{2}^{\rm obs}$. This has been verified by model study [44]. Thus, positive residual background remains in the linear fit intercept in the analysis of $\Delta\gamma$ as a function of $v_{2}^{\rm obs}$ [24], although the background is significantly reduced compared to that in the inclusive $\Delta\gamma$ measurement. Ideally, in the ES analysis, one would want to select events with the $v_{2}^{\rm obs}$ values of all resonances (background sources) to be zero, however, such selection is experimentally insurmountable.
@@
It has been recently proposed by Xu et al. [36] to select events according to the $\hat{q}_{2,\textsc{poi}}^{2}$ of the POIs, and study the $\Delta\gamma$ as a function of $v_{2}$ of the POIs. This is similar in spirit to the ES method above where the events are selected by $v_{2}^{\rm obs}$ and the $\Delta\gamma$ is examined as a function of the same $v_{2}^{\rm obs}$ variable. Here, the selection variable $\hat{q}_{2,\textsc{poi}}^{2}$ and the $v_{2}$ variable are different, although related as they are computed by the same POIs. The $\hat{q}_{2}^{2}$ variable is an event-by-event quantity connected to two-particle cumulant anisotropy (cf Eq. 5), whereas $v_{2}$ can be measured with respect to $\psi$ from another kinematic region like in the ES method. (Of course $v_{2}$ can also be computed by the two-particle cumulant method, in which case it is more directly connected to $\hat{q}_{2}^{2}$.) The variations in the event-by-event $\hat{q}_{2,\textsc{poi}}^{2}$ quantity, within a given narrow centrality bin, are mainly from statistical fluctuations. Since the $v_{2}$ in those $\hat{q}_{2,\textsc{poi}}^{2}$-selected events are related to $\hat{q}_{2,\textsc{poi}}^{2}$ itself, the variations in $v_{2}$ is also primarily of statistical nature. (Note, we simply use $v_{2}$ here, instead of the $\left\langle v_{2}\right\rangle$ for ESE, to indicate that the $v_{2}$ in ESS is of statistical nature although it is an “average” $v_{2}$ in events selected by $\hat{q}_{2,\textsc{poi}}^{2}$.) This technique selects on the statistical fluctuations in event shape, not on the dynamical fluctuations in $\left\langle v_{2}\right\rangle$ like in the ESE method, and it is referred to as the Event-Shape-Selection (ESS) method [36]. The most important difference in the ESS method from the ESE method is that the $\hat{q}_{2}^{2}$ is computed from the POIs in the former but in the later the phase-space where to compute $\hat{q}_{2}^{2}$ is displaced from that of the POIs’. In both approaches, one groups events (within a given narrow centrality bin) according to $\hat{q}_{2}^{2}$ and studies the $\Delta\gamma$ variable as a function of the $v_{2}$ of the POIs.
@@
Several combinations of $\hat{q}_{2,\textsc{poi}}^{2}$ and $v_{2}$ variables have been investigated in Ref. [36]. The $q_{2,\textsc{poi}}^{2}$ can be computed using single particle azimuthal angles of the POIs by Eq. 5. It can also be defined by pairs of POIs similar to Eq. 5; in this case, the $q_{2,{\rm pair}}^{2}$ of pairs is calculated by substituting the $i^{\rm th}$ particle’s $\phi_{i}$ by the azimuthal angle of the $i^{\rm th}$ pair in Eq. 5, and the sum runs over all pairs of POIs in the event. The azimuthal angle of a pair is defined to be that of the total momentum of the two POIs (vector sum of their momenta). The normalized $\hat{q}_{2,{\rm pair}}^{2}$ is then